OpenStudy (anonymous):

solve for y (y-7)/(y-1)=[(y-6)/(y-3)]+1

OpenStudy (anonymous):

y = -3 or y =4

OpenStudy (anonymous):

dam its not 6 and -2

OpenStudy (dumbcow):

\[\frac{y-7}{y-1} = \frac{y-6}{y-3}+1\] combine right side into single fraction \[\frac{y-7}{y-1} = \frac{y-6}{y-3}+\frac{y-3}{y-3} = \frac{2y-9}{y-3}\] cross-multiply \[(y-7)(y-3) = (2y-9)(y-1)\] \[y^{2}-10y+21 = 2y^{2}-11y+9\] combine like terms, set equal to zero \[y^{2}-y-12 = 0\] Factor \[(y-4)(y+3)=0\] y = 4, y = -3